# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 3.1.2.  Singular perturbation theory
Model reduction by SPT is a nonlinear subspace method based on time-scale separation of a FOM [80]. Assume that the FOM (2) is originally in “singular perturbation canonical form” [27]:
@@
where $\varepsilon\ll 1$ and $\bm{x}_{1}(t)\in\mathbb{R}^{n_{z}}$ and $\bm{x}_{2}(t)\in\mathbb{R}^{n_{x}-n_{z}}$. Equations 12a and 12b collect the “slow” and “fast dynamics”, respectively. We assume that exponential stability of the fast dynamics holds uniformly.
The key idea of perturbation theory is that the system response does not change significantly when $\varepsilon$ is perturbed to even smaller values. When setting $\varepsilon=0$ in Eq. 12, the fast dynamics degenerate into quasi-stationary equations, yielding a ROM defined by:
@@
Assuming that Eq. 13 is index-one, the corresponding slow manifold is given by:
=$\mathcal{M}=\{(\bm{x}_{1},\bm{x}_{2},\bm{u})\in\mathcal{\mathbb{R}}^{n_{z}}\times\mathbb{R}^{n_{x}-n_{z}}\times\mathbb{R}^{n_{u}}:\bm{f}_{2}(\bm{x}_{1},\bm{x}_{2},\bm{u})=\bm{0}\}\,.$ (14)
Typically, a process model is initially not in the form (12) and identifying the perturbation parameter $\varepsilon$ is not obvious [11]. Often, we seek for a linear transformation, $\bm{x}(t)=\mathbf{T}_{1}\bm{x}_{1}(t)+\mathbf{T}_{2}\bm{x}_{2}(t)$. For chemical systems, $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$ are sometimes determined based on a time scale assumption on energy and mass transport [81, 82]. More rigorous approaches search for a subset of $\mathcal{X}$ or $\mathcal{X}\times\mathcal{U}$ whose defect of invariance is sufficiently small [27, 83, 84].
An important property of Eq. 13 is that the steady-state response remains exact. However, the degeneration of the fast dynamics occasionally introduces undesirable dynamic artifacts, especially a non-physical inverse response (IR) [85, 86]. The appearance of an IR depends on the degree of reduction and the choice of $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$ [87]. In control, a sign-reversed model gain can degrade the closed-loop behavior of NMPC toward infeasibility or instability. However, a false IR is only problematic when appearing on the prediction time grid of the ROM application, e.g., NMPC, where an extremely fast IR may be uncritical.
3.1.3.  Redisualization method
Closely related to the singular perturbation method is the “POD residualization” (POD-Res) or “nonlinear-POD-Galerkin” method [88, 13]. Instead of Galerkin projection onto a linear subspace via POD truncation, a nonlinear subspace ROM is obtained by considering the limit of infinitely fast (degenerate or “slaved”) dynamics. Let the POD modes $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2},\mathbf{U}_{3}]$ so that $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$ and $\mathbf{U}_{2}\in\mathbb{R}^{n_{x}\times n_{q}}$, where $1<n_{q}\leq n_{x}-n_{z}$. Accordingly, we first extend Eq. 11 to:
@@ 3.1.3.  Redisualization method
Then, by residualization of Eq. 15b, i.e., setting $\dot{\bm{z}}_{2}(t)=\bm{0}$, we obtain [17]:
@@
Despite the POD-based partitioning of states in the first step, the residualization provides a nonlinear subspace due to Eq. 16b. In [60], the authors compared POD-Galerkin and POD-Res, finding that low-order POD-Res ROMs are more accurate than POD-Galerkin ROMs but also more computationally expensive without hyperreduction.
@@ 3.1.4.  Manifold-Galerkin method
Applying machine learning (ML) to approximate a manifold from snapshot data is called “manifold learning” (MFL) or “representation learning” [91]. MFL encompasses a wide array of tools and techniques that can be grouped into (i) ANN approaches, including autoencoder networks (AENs) [92, 93] and self-organizing maps [94], (ii) spectral methods, such as classical PCA or kernel PCA [95, 96], locally linear embeddings [97], Laplacian eigenmaps [98], and diffusion maps [99], and (iii) probabilistic methods [100]. These methods can also be combined, e.g., AENs and classical PCA [101].
Not all MFL methods are suitable for MOR via Galerkin projection. The subset of viable options includes AENs, self-organizing maps, and kernel PCA [36]. Herein, we focus on MOR using AENs. In particular, the simultaneous strength and weakness of AENs is their feedforward structure, which renders AENs conceptually and numerically simple to deal with. However, although AENs can be successfully applied in many cases, some manifolds cannot be globally represented by a single parameterization as provided by an AEN.
@@
The structure of an AEN is illustrated by Fig. 2. We here consider “undercomplete” AENs, where $n_{z}<n_{x}$. First, the “encoder” network $\bm{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{n_{z}}$, $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ open, projects the high-dimensional input vector $\bm{x}\in\mathcal{X}$ to a lower-dimensional representation $\bm{z}\in\mathbb{R}^{n_{z}}$. Often, $\bm{z}$ are called “features” or “latent variables”. In the second step, the “decoder” network $\bm{\phi}^{\dagger}:\mathcal{Z}\rightarrow\mathbb{R}^{n_{x}}$, where $\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}$, reconstructs the high-dimensional vector from the features. The images of the two mappings are $\bm{\phi}(\mathcal{X})\subseteq\mathcal{Z}$ and $\bm{\phi}^{\dagger}(\mathcal{Z})\subseteq\mathcal{X}$. The standard unsupervised AEN training problem minimizes the mean squared reconstruction error (MSE), $\sum_{k=1}^{N}\lVert\bm{x}^{(k)}-\bm{\phi}^{\dagger}\circ\bm{\phi}^{\dagger}(\bm{x}^{(k)})\rVert_{2}^{2}$, given a data set $\mathcal{D}=\{\bm{x}^{(k)}\}_{k=1}^{N}$ [102]. Assuming that $\bm{\phi}^{\dagger}$ is a one-to-one immersion, the subset $\mathcal{M}_{\mathcal{X}}\triangleq\{\bm{\phi}^{\dagger}(\hat{\bm{z}}):\hat{\bm{z}}\in\mathcal{Z}\}$ is an $n_{z}$-dimensional embedded submanifold of $\mathcal{X}$. When using an MSE training loss, the composition $\bm{\phi}^{\dagger}\circ\bm{\phi}$ approximates an orthogonal projection from $\mathcal{X}$ to $\mathcal{M}_{\mathcal{X}}$.
@@
Intrusive MOR based on AENs is presented in [103, 36] and refined in [101, 104]. Here, we adopt the method of [36] and we comment on the refinements below. Training an AEN on snapshots approximates the slow manifold by $\mathcal{M}_{\mathcal{X}}$. Differentiation of the approximation $\bm{x}(t)\approx\bm{\phi}^{\dagger}(\bm{z}(t))$ and insertion into Eq. 2 yields:
$\begin{split}\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))\dot{\bm{z}}(t)&=\bm{f}(\bm{\phi}^{\dagger}(\bm{z}(t)),\bm{u}(t))\,,\\ \bm{x}_{\mathrm{r}}(t)&=\bm{\phi}^{\dagger}(\bm{z}(t))\,,\end{split}$ (17)
@@
where $\dot{\bm{z}}(t)\in\mathbb{R}^{n_{z}}$ and $\bm{x}_{\mathrm{r}}(t)\in\mathbb{R}^{n_{x}}$. Clearly, the map $\bm{\phi}^{\dagger}$ (and thus the activation function) need to be differentiable. To obtain a well-posed ROM, we multiply Eq. 17 by $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{\mathrm{T}}$:
@@
As a result, we obtain a “manifold-Galerkin” projection of FOM (2) onto $\mathcal{M}_{\mathcal{X}}$ [36]. By applying the definition of the pseudoinverse, $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{+}$, Eq. 18 can be finally brought into the form of Eq. 4:
@@
Note that Eq. 19 does not involve the encoding $\bm{\phi}$. While $\bm{\phi}^{\dagger}(\bm{z}(t))$ is an embedding of $\mathcal{Z}$ into $\mathcal{X}$, the map $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{+}$ is a linear projection from $\mathbb{R}^{n_{x}}$ to $\mathbb{R}^{n_{z}}$ (more precisely from the tangent space to $\mathcal{X}$ at $\bm{\phi}^{\dagger}(\bm{z}(t))$ onto the tangent space to $\mathcal{Z}$ at $\bm{z}(t)$). When using a linear AEN, then Eq. 19 reduces to Eq. 11. In practice, we employ Eq. 18 rather than Eq. 19 to circumvent matrix inversion or SVD associated with pseudoinversion in the model. To improve ROM accuracy, the standard AEN training problem may be extended by an additional velocity loss [104]. We regard such an approach as a step towards “physics-informed” model reduction.
There exist alternative options to using manifold-Galerkin projection and the pseudoinverse. One alternative approach, that is also suitable when applying non-orthogonal (oblique) projections [104], is to apply the left-inverse property, $\mathrm{D}\bm{\phi}(\bm{\phi}^{\dagger}(\bm{z}))\circ\mathrm{D}\bm{\phi}^{\dagger}(\bm{z})=\mathbf{I}$, to Eq. 17. However, within preliminary studies, we found that this approach is sensitive to projection errors and therefore requires an advanced AEN training strategy.
Fig. 2: Illustration of an autoencoder network.
The manifold-Galerkin projection using AENs increases the number of FLOPS considerably due to the added decoder network and its Jacobian. On the positive side, the slow manifold is represented by a feedforward ANN rather than by a set of implicit equations. For spatially distributed systems, computational improvements can be achieved by using convolutional instead of fully connected AENs [105, 106].
@@
We are particularly interested in systems with input. Although [103, 36, 104] consider input-affine or parametric systems, these works do not account for the effect of inputs or parameters on the slow manifold, cf. Eq. 6. Instead, a trivial bundle $\mathcal{M}=\mathcal{M}_{\mathcal{X}}\times\mathcal{U}$ is constructed. However, we expect this simplification to demand an unnecessarily high-dimensional slow manifold. Consequently, we employ the concepts from Section 2 to propose a respective extension below.
@@ 3.1.5.  Manifold-Galerkin method with inputs
We extend the manifold-Galerkin method [36] to account for the effect of system inputs on the slow manifold $\mathcal{M}$. Restricting ourselves to AENs, we are generally interested in maps of the form $\bm{\varphi}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{Z}\times\mathcal{U}$ and $\bm{\varphi}^{\dagger}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathcal{X}\times\mathcal{U}$. The corresponding manifold is =$\mathcal{M}=\{\bm{\varphi}^{\dagger}(\hat{\bm{z}},\hat{\bm{u}}):\hat{\bm{z}}\in\mathcal{Z},\hat{\bm{u}}\in\mathcal{U}\}\subseteq\mathcal{X}\times\mathcal{U}$, where $\bm{\varphi}^{\dagger}\circ\bm{\varphi}(\mathcal{X}\times\mathcal{U})=\mathcal{M}$ and $\bm{\varphi}\circ\bm{\varphi}^{\dagger}=\mathbf{id}$ (identity). For convenience, we substructure $\bm{\varphi}^{\dagger}$ by stacking the maps $\bm{\psi}^{\dagger}:\mathcal{Z}\times\mathcal{U}\rightarrow\mathcal{X}$ and $\mathbf{id}:\mathcal{U}\rightarrow\mathcal{U}$, so that $\mathcal{M}_{\mathcal{X}}(\bm{u}_{0})=\{\bm{\psi}^{\dagger}(\hat{\bm{z}},\bm{u}_{0}):\hat{\bm{z}}\in\mathcal{Z}\}$. Similarly, we substructure $\bm{\varphi}$ into the stacked maps $\bm{\psi}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{Z}$ and $\mathbf{id}:\mathcal{U}\rightarrow\mathcal{U}$.
Based on Eq. 4, we define pointwise:
$\bm{h}(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}})\triangleq\hat{\bm{x}}-\bm{\psi}^{\dagger}(\hat{\bm{z}},\hat{\bm{u}}),$ (20)
so that Eq. 7 writes equivalently:
$\forall(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}})\in\mathcal{M}:\;\bm{f}(\hat{\bm{x}},\hat{\bm{u}})-\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\hat{\bm{z}},\hat{\bm{u}})\bm{f}_{\mathrm{r}}(\hat{\bm{z}},\hat{\bm{u}})=\bm{0},$ (21)
where we use $\mathrm{D}_{\bm{u}}\bm{\psi}^{\dagger}(\,\cdot\,)\equiv\mathbf{0}$ as the derived case of $\mathrm{D}_{\bm{x}}\bm{h}(\,\cdot\,)^{-1}\mathrm{D}_{\bm{u}}\bm{h}(\,\cdot\,)\equiv\mathbf{0}$ (strict causality). As a consequence, we can redefine $\bm{\psi}$ by $\bm{\psi}:\mathcal{X}\rightarrow\mathcal{Z}$ so that the encoding only depends on $\bm{x}(t)$. This structural restriction guarantees that $\hat{\bm{z}}$ is not algebraically coupled to $\hat{\bm{u}}$. Consequently, we will apply an AEN training that minimizes the loss $\sum_{k}\lVert\bm{x}^{(k)}-\bm{\psi}^{\dagger}(\bm{\psi}(\bm{x}^{(k)}),\bm{u}^{(k)})\rVert_{2}^{2}$. We call our approach “manifold learning with inputs” (MFLu). Next, we project $\bm{f}$ onto $\mathcal{M}$:
$\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t))\,\dot{\bm{z}}(t)+\underbrace{\mathrm{D}_{\bm{u}}\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t))}_{=\,\mathbf{0}\text{ (see above)}}\,\dot{\bm{u}}(t)=\bm{f}(\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t)),\bm{u}(t)).$ (22)
Finally, after left-multiplication by the pseudoinverse $\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t))^{+}$, or alternatively by $\mathrm{D}_{\bm{z}}\bm{\psi}(\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t)))$, we obtain the ROM:
@@
This extension is related to studies on parametric model reduction; see [77, 107] for an overview. However, present works mostly deal with stationary parametric systems and use extensions of linear basis methods. In particular, [108, 109] combine a reduced POD basis and polynomial interpolation to build low-dimensional steady-state models of parametric problems. Refinements of this approach are found in [110, 111]. Reduction of parametric systems via manifold learning has also been investigated in the literature: To reduce the dimensionality of steady parametric flow problems, [112] present a strategy based on isomaps, and [113] apply kernel PCA. In [114], diffusion maps are used to learn submanifolds of the input-output space for reducing the parameter dimension. However, these works do not consider MOR of non-autonomous dynamical systems.

### v2

@@ 3.1.2. Singular perturbation theory
Model reduction by SPT is a nonlinear subspace method based on time-scale separation of a FOM [145, 67]. Assume that the FOM (2) is originally in singular perturbation canonical form [140]:
@@
where $\varepsilon\ll 1$ and $\bm{x}_{1}(t)\in\mathbb{R}^{n_{z}}$ and $\bm{x}_{2}(t)\in\mathbb{R}^{n_{x}-n_{z}}$. Eqs. (13a) and (13b) collect the slow and fast dynamics, respectively. We assume that exponential stability of the fast dynamics holds uniformly.
The key idea of perturbation theory is that the system response does not change significantly when $\varepsilon$ is perturbed. When setting $\varepsilon=0$ in (13), the fast dynamics degenerate into quasi-stationary equations, yielding a ROM defined by:
@@
Assuming that (14) is index-one, the corresponding slow manifold is given by:
$\mathcal{M}=\{(\hat{\bm{x}}_{1},\hat{\bm{x}}_{2},\hat{\bm{u}})\in\mathcal{\mathbb{R}}^{n_{z}}\times\mathbb{R}^{n_{x}-n_{z}}\times\mathbb{R}^{n_{u}}:\bm{f}_{2}(\hat{\bm{x}}_{1},\hat{\bm{x}}_{2},\hat{\bm{u}})=\bm{0}\}\,.$ (15)
Overall, the projection of (13) onto $\mathcal{M}$ is inherently defined via the degeneration.
Typically, a process model is initially not in the form (13). Identifying the perturbation parameter $\varepsilon$ is often not obvious and frequently based on human assumptions rather than system-theoretic arguments. In many cases, we seek a linear transformation, $\bm{x}(t)=\mathbf{T}_{1}\bm{x}_{1}(t)+\mathbf{T}_{2}\bm{x}_{2}(t)$. For chemical process systems, $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$ are sometimes determined based on a time scale assumption on energy and mass transport [293, 244]. More rigorous approaches search for a subset of $\mathcal{X}$ or $\mathcal{X}\times\mathcal{U}$ whose defect of invariance is sufficiently small [140, 109, 130]. Further, note that implicit condensation [121, 202] and static condensation [104] in mechanics are closely connected to SPT. For example, the needed slow/fast time-scale separation for nonlinear vibrations was examined in [104, 268].
An important property of (14) is that the steady-state response remains exact. However, the degeneration of the fast dynamics occasionally introduces undesirable dynamic artifacts, especially a non-physical inverse response [20, 234]. In other words, the degenerated finite-$\varepsilon$ dynamics have a substantial macroscopic effect when (13) does not constitute a proper normal form [242]. Also, the appearance of an inverse response depends on the degree of reduction and the determination of $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$ [125]. In control, a sign-reversed model gain can degrade the closed-loop behavior of a model-based controller toward infeasibility or instability. However, a false inverse response is only problematic when appearing on the prediction time grid of the ROM application, where an extremely fast inverse response may be uncritical.
3.1.3. Residualization method
Closely related to the singular perturbation method is the POD residualization (POD-Res) or nonlinear-POD-Galerkin method [70, 92]. Instead of Galerkin projection onto a linear subspace via POD truncation, a nonlinear subspace ROM is obtained by considering the limit of infinitely fast (degenerate or slaved) dynamics. Let the POD modes $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2},\mathbf{U}_{3}]$ so that $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$ and $\mathbf{U}_{2}\in\mathbb{R}^{n_{x}\times n_{q}}$, where $1<n_{q}\leq n_{x}-n_{z}$. Accordingly, we first extend (12) and truncate $\mathbf{U}_{3}$ so that:
@@ 3.1.3. Residualization method
Then, by residualization of (16b), i.e., setting $\dot{\bm{z}}_{2}(t)=\bm{0}$, we obtain [269]:
@@
Despite the POD-based partitioning of states in the first step, the residualization provides a nonlinear subspace due to (17b). In [256], the authors compared POD-Galerkin and POD-Res, finding that low-order POD-Res ROMs are more accurate than POD-Galerkin ROMs but also more computationally expensive without hyperreduction.
@@ 3.1.4. Manifold-Galerkin method
Applying machine learning to approximate a manifold from snapshot data is called manifold learning (MFL) or representation learning [179, 279]. MFL encompasses a wide array of tools and techniques that can be grouped into (i) learning methods, including autoencoder networks (AENs) [14, 150] and self-organizing maps [144], (ii) spectral methods, such as classical PCA or kernel PCA [258, 194], locally linear embeddings [247], Laplacian eigenmaps [19], and diffusion maps [52], and (iii) probabilistic methods [82]. These methods can also be combined, e.g., classical PCA and quadratic decoding [15, 78] or classical PCA and AENs [227].
@@
Not all MFL methods are suitable for MOR via Galerkin projection. The subset of viable options includes AENs, self-organizing maps, and kernel PCA [159]. Herein, we focus on MOR using AENs, which are a prominent choice, e.g., [136, 159, 85]. We here consider undercomplete AENs, where $n_{z}<n_{x}$. First, the encoder network $\bm{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{n_{z}}$, $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ open, projects the high-dimensional input vector $\bm{x}\in\mathcal{X}$ to a lower-dimensional representation $\bm{z}\in\mathbb{R}^{n_{z}}$. Often, $\bm{z}$ are called features or latent variables. In the second step, the decoder network $\bm{\phi}^{\dagger}:\mathcal{Z}\rightarrow\mathbb{R}^{n_{x}}$, where $\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}$, reconstructs the high-dimensional vector from the features. The images of the two mappings are $\bm{\phi}(\mathcal{X})\subseteq\mathcal{Z}$ and $\bm{\phi}^{\dagger}(\mathcal{Z})\subseteq\mathcal{X}$. The standard unsupervised AEN training problem minimizes the mean squared reconstruction error, $\sum_{k=1}^{N}\lVert\bm{x}(t_{k})-\bm{\phi}^{\dagger}\circ\bm{\phi}(\bm{x}(t_{k}))\rVert_{2}^{2}$, given a data set $\mathcal{D}=\{\bm{x}(t_{k})\}_{k=1}^{N}$ [286]. Assuming that $\bm{\phi}^{\dagger}$ is a one-to-one immersion, the subset $\mathcal{M}_{\mathcal{X}}\triangleq\{\bm{\phi}^{\dagger}(\hat{\bm{z}}):\hat{\bm{z}}\in\mathcal{Z}\}$ is an $n_{z}$-dimensional embedded submanifold of $\mathcal{X}$. When using an mean squared training loss, the composition $\bm{\phi}^{\dagger}\circ\bm{\phi}$ approximates an orthogonal projection from $\mathcal{X}$ to $\mathcal{M}_{\mathcal{X}}$.
@@
Intrusive MOR based on AENs is presented in [136, 159] and refined, e.g., in [227, 209]. Here, we adopt the method of [159] and we comment on the refinements below. Training an AEN on snapshots approximates the slow manifold by $\mathcal{M}_{\mathcal{X}}\times\mathcal{U}$. Differentiation of the approximation $\bm{x}(t)\approx\bm{\phi}^{\dagger}(\bm{z}(t))$ and insertion into (2) yields:
[…]
@@
where $\dot{\bm{z}}(t)\in\mathbb{R}^{n_{z}}$ and $\bm{x}_{\mathrm{r}}(t)\in\mathbb{R}^{n_{x}}$. Clearly, the map $\bm{\phi}^{\dagger}$ (and thus the activation function) need to be differentiable. To obtain a well-posed ROM, we multiply (18) by $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{\mathrm{T}}$:
@@
As a result, we obtain a manifold-Galerkin projection of FOM (2) onto $\mathcal{M}_{\mathcal{X}}$ [159]. By applying the definition of the pseudoinverse, $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{+}$, Eq. (19) can be finally brought into the form of (4):
@@
In practice, we employ (19) rather than (20) to circumvent matrix inversion or SVD associated with pseudoinversion in the model. Note that both (18) and (20) do not involve the encoding $\bm{\phi}$. While $\bm{\phi}^{\dagger}(\bm{z}(t))$ is an embedding of $\mathcal{Z}$ into $\mathcal{X}$, the map $\mathrm{D}\bm{\phi}^{\dagger}(\bm{z}(t))^{+}$ is a linear projection from $\mathbb{R}^{n_{x}}$ to $\mathbb{R}^{n_{z}}$ (more precisely from the tangent space to $\mathcal{X}$ at $\bm{\phi}^{\dagger}(\bm{z}(t))$ onto the tangent space to $\mathcal{Z}$ at $\bm{z}(t)$). When using a linear AEN, then (20) reduces to (12). Finally, we highlight the close connection between (20) and other approaches using a single parameterization/decoder map, e.g., [240, 130].
An alternative to pseudoinversion is to apply the left-inverse property, $\mathrm{D}\bm{\phi}(\bm{\phi}^{\dagger}(\,\cdot\,))\circ\mathrm{D}\bm{\phi}^{\dagger}(\,\cdot\,)=\mathbf{I}$, to (18). This approach is also valid for non-orthogonal (oblique) projections and yields [209]:
$\dot{\bm{z}}(t)=\mathrm{D}\bm{\phi}(\bm{\phi}^{\dagger}(\,\cdot\,))\bm{f}(\bm{\phi}^{\dagger}(\bm{z}(t)),\bm{u}(t))$ (21)
instead of (20a). However, computing $\mathrm{D}\bm{\phi}(\bm{\phi}^{\dagger}(\,\cdot\,))$ can be expensive and the precise satisfaction of the left-inverse property is a crucial prerequisite of this method.
Despite their intuitive structure and popularity in recent years, standard dense AENs suffer from limited applicability, poor scaling in the number of parameters and training costs, expensive model evaluation, moderate accuracy, and potential singularity issues. We discuss these challenges next. First, the simultaneous strength and weakness of AENs is their feedforward structure, which renders AENs conceptually and numerically simple to deal with. However, some manifolds cannot be globally represented by a single parameterization/decoder as provided by an AEN. Consequently, AENs may not be applicable to some problems. On the positive side, the slow manifold is represented by a feedforward structure rather than by a set of implicit equations, see (4b).
A second issue is that very high $n_{x}$ and $n_{z}$ cause an explosion in the parameters of dense encoder and decoder networks, resulting in intractable training and model evaluation. Additionally, embedding an AEN into a ROM increases the number of floating point operations of model evaluation considerably. For spatially distributed systems, computational improvements can be achieved by using convolutional instead of fully connected AENs [74, 246]. Moreover, there exist more parameter-efficient approaches to realize the autoencoder structure, e.g., polynomials with compressed tensor coefficients have linear complexity scaling [93, 279]. On the other hand, we remark that the FOM equations, $\bm{f}$ in (20), are often already computationally expensive, necessitating a hyperreduction step in any case. Then, the AEN evaluation costs are secondary as the right-hand side of (20a) will be ultimately replaced by a cheaper function.
To improve ROM accuracy, the standard AEN training problem may be extended by a robustness loss [308] as well as a velocity loss [209], which we regard as a step towards physics-informed model reduction. Furthermore, a standard AEN often faces injection errors, i.e., violating the identity $\bm{\varphi}\circ\bm{\varphi}^{\dagger}=\mathbf{id}$ or equivalently $\bm{\varphi}^{\dagger}\circ\bm{\varphi}\circ\bm{\varphi}^{\dagger}\circ\bm{\varphi}=\bm{\varphi}^{\dagger}\circ\bm{\varphi}$ (idempotence). This identity can be included into the training problem [279] or enforced structurally [209]. Moreover, learning an invariant foliation instead of a single manifold may increase accuracy of the reduced space [278, 279].
The AEN-based manifold-Galerkin method is sensitive to ill-conditioning of the Jacobian $\mathrm{D}\bm{\phi}^{\dagger}(\,\cdot\,)$. Unfortunately, the Jacobian from a standard AEN training may be arbitrarily ill-conditioned. However, conditioning issues can be addressed by including a rank loss [200] in the training or employing an invertible network structure possessing an exact (full-rank) Jacobian left-inverse [209].
@@
Herein, we are particularly interested in systems with input. Although [159, 209] consider input-affine or parametric systems, these works do not account for the effect of inputs or parameters on the slow manifold, cf. (7). Instead, a trivial bundle $\mathcal{M}=\mathcal{M}_{\mathcal{X}}\times\mathcal{U}$ is constructed. However, we expect this simplification to demand an unnecessarily high-dimensional slow manifold for some systems. Consequently, we employ the concepts from Section 2 to propose a respective extension below.
@@ 3.1.5. Manifold-Galerkin method with inputs
We extend the machine learning manifold-Galerkin method [159] to account for the parametric effect of system inputs on the slow manifold $\mathcal{M}$ and $\mathcal{M}_{\mathcal{X}}$. This explicit parametric dependency has not been accounted for in the learning strategy proposed in [159] and related works (see Sections 3.1.4 and 3.2.4).
[…]
Given the above characterization of $\mathcal{M}$, Eq. (8) becomes:
$\forall(\hat{\bm{x}},\hat{\bm{z}},\hat{\bm{u}})\in\mathcal{M}:\;\bm{f}(\hat{\bm{x}},\hat{\bm{u}})-\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\hat{\bm{z}},\hat{\bm{u}})\bm{f}_{\mathrm{r}}(\hat{\bm{z}},\hat{\bm{u}})=\bm{0}{.}$ (22)
To guarantee that $\bm{z}(t)$ is not algebraically coupled to $\bm{u}(t)$, we simplify $\bm{\psi}$ to $\bm{\psi}:\mathcal{X}\rightarrow\mathcal{Z}$. Consequently, the encoding only depends on $\bm{x}(t)$. We apply an AEN training that minimizes the loss $\sum_{k=1}^{N}\lVert\bm{x}(t_{k})-\bm{\psi}^{\dagger}(\bm{\psi}(\bm{x}(t_{k})),\bm{u}(t_{k})\rVert_{2}^{2}$. Similarly to Section 3.1.4, this loss may be extended to incorporate system knowledge and reduce projection errors. We call our approach manifold learning with inputs (MFLu). Next, we project $\bm{f}$ onto $\mathcal{M}_{\mathcal{X}}(\bm{u}(t))$:
$\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t))\,\dot{\bm{z}}(t)\bm{f}(\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t)),\bm{u}(t)).$ (23)
Finally, after left-multiplication by the pseudoinverse $\mathrm{D}_{\bm{z}}\bm{\psi}^{\dagger}(\bm{z}(t),\bm{u}(t))^{+}$, we obtain the ROM:
@@
This extension is related to studies on parametric model reduction; see [22, 24] for an overview. However, present works mostly deal with stationary parametric systems and use extensions of linear basis methods. In particular, [177, 38] combine a reduced POD basis and polynomial interpolation to build low-dimensional steady-state models of parametric problems. Refinements of this approach are found in [182, 118]. Reduction of parametric systems via manifold learning has also been investigated in the literature: To reduce the dimensionality of steady parametric flow problems, [72] present a strategy based on isomaps, and [57] apply kernel PCA. In [120], diffusion maps are used to learn submanifolds of the input-output space for reducing the parameter dimension. However, these works do not consider MOR of non-autonomous dynamical systems with inputs. Finally, we remark that MFLu is conceptually not restricted to AENs and may be adapted to other types of manifold learning.